Average movie prices in the United States are, in general, lower than in other countries. It would cost $81.22 to buy three tickets in Japan plus two tickets in Switzerland. Three tickets in Switzerland plus two tickets in Japan would cost $77.18 . How much does an average movie ticket cost in each of these countries?
Japan the average cost is $___
Switzerland the average cost is $___

Answer by ankor@dixie-net.com(22740)

It would cost $81.22 to buy three tickets in Japan plus two tickets in Switzerland.
Three tickets in Switzerland plus two tickets in Japan would cost $77.18 .
How much does an average movie ticket cost in each of these countries?
:
let s = the cost of ticket in Switzerland
let j = the cost in Japan
:
Write an equation for each transaction, we will use elimination here.
:
3j + 2s = 81.22
2j + 3s = 77.18
multiply the 1st eq by 3, and the 2nd eq by 2
9j + 6s = 243.66
4j + 6s = 154.36
-------------------
5j + 0 = 89.30
j = 89.30/5
j = $17.86, the cost of a ticket in Japan
:
Use the 1st equation to find s, replace j with 89.30
3(17.86) + 2s = 81.22
53.58 + 2s = 81.22
2s = 81.22 - 53.58
2s = 27.64
s = 27.64/2
s = $13.82 the cost of a ticket in Switzerland
:
;
Use the 2nd equation; 2j + 3s = 77.18, to check our answers
2(17.86) + 3(13.82) = 77.18
35.72 + 41.46 = 77.18
7718 = 77.18, confirms our solutions
:
;
Japan the average cost is $_17.86__
Switzerland the average cost is $_13.82__